Last year's dividend (D0) was$3.20per share. The constant growth in dividends=6.25%. Required return=12%.

Requirement 1:

What is the current price? (Do not round intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).)

  Current price $   
Requirement 2:

What will the price be four years from now? sixteen years? Hint: recall the price grows at the constant growth rate of dividends. (Do not round intermediate calculations. Round your answers to 2 decimal places (e.g., 32.16).)

  Four years $   
  Sixteen years $   
Answer #1

Solution:

D0 =$3.20

Growth rate (g)= 6.25%

D1= 3.20*(1+0.0625)= $3.4

Required return ( r) = 12%

Current price ( P0) = D1/(r- g)

P0 = 3.40/(0.12 - 0.0625)

P0 = $59.13

Current price = $59.13

Price be four years from now ( P4) = 59.13*(1.0625)^4= $75.36

Price be 16 years from now ( P16) = 59.13*(1.0625)^16= $155.98